OpenStudy (anonymous):

write the answer in slope intercept and standard form. parallel to y=-5 and through (2,7)

OpenStudy (anonymous):

y-7=-5(2)? or would i have to find the slope first to say: y-7=slope(x-2)?

OpenStudy (anonymous):

it would have to obviously be 69

sam (.sam.):

If the two lines are parallel, what will be the same?

OpenStudy (anonymous):

what do you mean? One of the variables (would slope and the y intercept be the same?)

sam (.sam.):

Yes the slope will be the same

sam (.sam.):

y=-5 is 

OpenStudy (anonymous):

now do i have to solve for slope?

sam (.sam.):

Point (2,7) would be in 

sam (.sam.):

Alright good luck :)

OpenStudy (anonymous):

what does this graph mean?